solve by graphing
7x-6y=30
6y-7y=-30

Answer by stanbon(75887)

pipe A cn fill a cistern in 6hrs & pipe B cn fill it in 8hrs. both pipes r opened simultaneously bt after 2hrs pipeA is closed. how many hrs willB take to fill the remaining partbof cistern
---------------------------
A time = 6 hrs/job ; rate = 1/6 job/hr
-------
B time = 8 hrs/job ; rate = 1/8 job/hr
-------
Together rate = 1/6 + 1/8 = (14/48) = 7/24 job/hr
-----------------------------------------
Note: work done = rate*time
-----------------------------------------------------
Equation:
work + work = 1 job
2(7/24) + (1/8)x = 1
14 + 3x = 24
3x = 14
x = 4 1/3 hrs (time for "B" to finish the job)
=================
Cheers,
Stan H.
==================